OpenStudy (anonymous):

Evaluate the integral: int/ 7u^2-(2/u)

OpenStudy (lgbasallote):

\[\int 7u^2 - \frac 2u?\]

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

split the integral.. \[\int 7u^2 du - \int \frac{2}{u} du\] do you know how to integrate those?

OpenStudy (anonymous):

no i do not

OpenStudy (lgbasallote):

what if i take out the constants.. \[7\int u^2 du - 2\int \frac 1u du\]

OpenStudy (lgbasallote):

do those ring a bell?

OpenStudy (anonymous):

5(1/u)^2?

OpenStudy (turingtest):

\[\int x^ndx={x^{n+1}\over n+1}+C\]and\[\int\frac1xdx=\ln|x|+C\]

OpenStudy (anonymous):

where are the n's coming from

OpenStudy (lgbasallote):

these are definitions and rules

OpenStudy (lgbasallote):

n is just a variable

OpenStudy (anonymous):

so it would be 7^(u+1)/u+1 and ln lul

OpenStudy (lgbasallote):

hmm nope..the base is supposed to be u

OpenStudy (lgbasallote):

\[\int u^3 du = \frac{u^{3+1}}{3+1}\] \[\int u^4 du = \frac{u^{4+1}}{4+1}\] getting the thread?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so what's u^2

OpenStudy (anonymous):

so is it 7(u^3/2+1)?

OpenStudy (anonymous):

9? or 4?

OpenStudy (lgbasallote):

what do you mean 9 or 4?

OpenStudy (anonymous):

okay i guess i dont know what u^2 is

OpenStudy (lgbasallote):

you were close! simplify \[7(\frac{u^3}{2+1})\]

OpenStudy (anonymous):

7u^3/21?

OpenStudy (lgbasallote):

what's 2+1?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

int udv=uv - int v du

OpenStudy (lgbasallote):

then it\s\[\frac{7u^3}{3}\] got it/

OpenStudy (anonymous):

yes i just shouldnt have multiplied the 2+1 times 7

OpenStudy (anonymous):

its saying that that isnt the correct answer though

OpenStudy (lgbasallote):

lol there's still \(2\int \frac 1udu\) remember

OpenStudy (anonymous):

arg.. i hate this

OpenStudy (lgbasallote):

lol look at what turing said \[\int \frac 1x dx = \ln x\] so what's \[\int \frac 1u du/\]

OpenStudy (anonymous):

ln u

OpenStudy (anonymous):

but i dont know where to put that now..

OpenStudy (lgbasallote):

so combine...

OpenStudy (lgbasallote):

\[\int 7u^2du = \frac{7u^3}{3}\] \[2\int \frac 1u du = 2\ln u\] so what's \[\int u^2 du - 2\int \frac 1u du?\]

OpenStudy (lgbasallote):

there's a 7 on the first term

OpenStudy (anonymous):

5(1/u)?

OpenStudy (anonymous):

or just 5? im not sure

OpenStudy (lgbasallote):

what? what did you do?

OpenStudy (anonymous):

i dont know im guessing because i dont know what to do

OpenStudy (anonymous):

14(1/u)?

OpenStudy (lgbasallote):

analyze my last post carefully

OpenStudy (lgbasallote):

you\'re just going to substitute

OpenStudy (anonymous):

7(1/u)^2?

OpenStudy (lgbasallote):

the question is \[\int 7u^2 du - \int \frac{2}{u} du\] correct?

OpenStudy (anonymous):

yes. i just havent learned about subtracting them before

OpenStudy (lgbasallote):

what's \[\int 7u^2 du?\]

OpenStudy (anonymous):

21u^3?

OpenStudy (lgbasallote):

how?

OpenStudy (anonymous):

i dont know im basically guessing because i do not know what to do!

OpenStudy (anonymous):

14u^3?

OpenStudy (anonymous):

7*(2/u)^2??

OpenStudy (lgbasallote):

but i already told you how to do it..why guess

OpenStudy (lgbasallote):

scroll up

OpenStudy (anonymous):

becuase im not understanding how your telling me to do it! it isnt making any sense to me!

OpenStudy (anonymous):

im really not good at this!

OpenStudy (lgbasallote):

do you agree that \[\int u^2du\] is in the form \[\int x^n dx?\]

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

\[\int x^n dx = \frac{x^{n+1}}{n+1}\]

OpenStudy (lgbasallote):

this is a rule

OpenStudy (lgbasallote):

this is called the power rule

OpenStudy (anonymous):

alright so what does that do for my problem

OpenStudy (lgbasallote):

in \[\int u^2 du \] which is the counterpart of x?

OpenStudy (anonymous):

u?

OpenStudy (lgbasallote):

correct and hwhat is the counterpart of n

OpenStudy (anonymous):

2

OpenStudy (turingtest):

side note (sorry to interrupt) @drumjockey83 memorize the part that says "common integrals" http://tutorial.math.lamar.edu/pdf/Calculus_Cheat_Sheet_All.pdf it will make your life so much easier

OpenStudy (anonymous):

thank you

OpenStudy (lgbasallote):

so how what will be \[\frac{x^{n+1}}{n+1}\] here?

OpenStudy (anonymous):

u^(2+1)/2+1?

OpenStudy (lgbasallote):

so what will be*

OpenStudy (lgbasallote):

correct

OpenStudy (lgbasallote):

so \[\int u^2 du \implies \frac{u^{2+1}}{2+1} \implies \frac{u^3}{3}\] do you get that?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so wjat os \[7\int u^2 du?\]

OpenStudy (anonymous):

7u^(2+1)/2+1?

OpenStudy (lgbasallote):

correct. but simplify

OpenStudy (anonymous):

7(u^3/3)?

OpenStudy (lgbasallote):

nnice

OpenStudy (anonymous):

thats the answer?

OpenStudy (lgbasallote):

nope..that's just for \[\int 7u^2 du\]

OpenStudy (anonymous):

okay i thought we already went over that

OpenStudy (lgbasallote):

now we solve \[2 \int \frac 1u du\] well im trying to go from the beginning because you're geting confused somewhere..idk which

OpenStudy (lgbasallote):

what would be the answer to that one

OpenStudy (anonymous):

2(1/u)?

OpenStudy (lgbasallote):

nope. here's a hint.another integration rule \[\int \frac 1x dx = \ln x\] so what's \[\int \frac 1u du?\]

OpenStudy (anonymous):

ln u

OpenStudy (lgbasallote):

so what's \[2\int \frac 1u du\]

OpenStudy (anonymous):

im not sure!! i dont know whay to do with that 2!!

OpenStudy (anonymous):

2 ln u???

OpenStudy (lgbasallote):

correct

OpenStudy (anonymous):

so whats the complete answer ?

OpenStudy (lgbasallote):

so what's \[\int 7u^2 du - \int \frac 2u du?\]

OpenStudy (anonymous):

7(u^3/3)-2 ln u

OpenStudy (lgbasallote):

correct

OpenStudy (lgbasallote):

congrats ^_^

OpenStudy (anonymous):

its saying that that isnt right

OpenStudy (lgbasallote):

do you have choices?